OpenStudy (anonymous):

If a dairyman has 300 pounds of milk testing 3% butterfat, how many pounds of skimmed milk (with no butterfat) must he remove to have milk testing 3.6% butterfat?

OpenStudy (anonymous):

hhhheeeelllppppp!!!!

OpenStudy (king):

1.2?

OpenStudy (anonymous):

lol can u show me the steps?

OpenStudy (ar43r):

imho 18 pounds

OpenStudy (anonymous):

i think its 1.8

OpenStudy (king):

hmm.ya 1.8 see 3% of 300=9 3.6% of 300=10.8 so if it has to increase skimmed milk must decrease by 10.8-9=1.8 pnds

OpenStudy (anonymous):

i like your logic :D

OpenStudy (dumbcow):

No the answer is 50 lbs If you only remove 1.8 then you are left with 298.2 298.2*.036 = 10.73 which is much higher than 9 since the amount of butterfat remains the same 300(.03) = (300-x)(.036) 9 = 10.8 -.036x .036x = 1.8 x = 50

OpenStudy (ar43r):

nicely done. have to note your solution.

OpenStudy (dumbcow):

thanks another way of looking at it 9/300 = .03 so 9/x = .036 --> x=250

OpenStudy (ar43r):

correct!